OpenStudy (yuii):

The function q(w)=3+5(w−1) represents the number of quarters in a bowl on week w. What does the value 5 represent in this situation? A. Five quarters are added to the bowl every week. B. The value of the quarters in the bowl on Week 1 was $5. C. Quarters were added to the bowl for 5 weeks. D. There were 5 quarters in the bowl on Week 1.

satellite73 (satellite73):

how many quarters were in the bowl in week 3? I.e. what is q(3)?

satellite73 (satellite73):

is it not clear? if not, say so and i will show you

OpenStudy (yuii):

I'm confused.

satellite73 (satellite73):

to find \(q(3)\) replace the w in \[q(w)=3+5(w−1)\] by 3 and compute \[q(3)=3+5(3-1)=?\]

OpenStudy (yuii):

q = 13/3

satellite73 (satellite73):

i don't see any fractions there

satellite73 (satellite73):

oh lord ok i see what happened, lets back up a second

satellite73 (satellite73):

it is "q OF 3" not "q times 3" q is not a variable, you are not solving for it

satellite73 (satellite73):

\[q(w)\] is a function , it is \[q(w)=3+5(w-1)\] and it is a number if you replace w by a number so for example \[q(3)=3+5(3-1)=13\]

satellite73 (satellite73):

it is not a letter to solve for i think that is how you got \(q=\frac{13}{3}\) right?

OpenStudy (yuii):

Yes.

satellite73 (satellite73):

ok is it a bit more clear now that \[q(3)=13\]?

OpenStudy (yuii):

Yes.

satellite73 (satellite73):

ok good, in this scenario, that means in week 3 there are 13 quarters in the bowl now try \[q(4)=3+5(4-1)=?\]

satellite73 (satellite73):

did you get a number?